anand123:

Suppose two drugs are routinely used for the treatment of thyroid dysfunction. Drug X is known to cure the disorder 75% of the time and costs $88. Drug Y is known to cure the disorder 60% of the time and costs $75. The two drugs work independent of each other. The two treatment plans are as follows: Plan A: Treatment with Drug X—if not effective, treatment with Drug Y Plan B: Treatment with Drug Y—if not effective, treatment with Drug X

SmokeyBrown:

Hi, and welcome to QuestionCove! I'm not sure exactly what the question, and I'll ask you to provide more context to the information, particularly what we're supposed to be answering. Still, I can break down the treatment plans by their cost and effectiveness, based on the information given. Plan A has a 75% chance of being effective with only Drug X, which costs $88. The remaining 25% of the time, the cost will be 88+75 = $163 with both drugs Plan B has a 60% chance of being effective with only Drug Y, which costs $75. The remaining 40% of the time, the cost will be $163 with both drugs Plan A has an effectiveness of 0.75 + (0.25*0.60) = 0.9 or 90% Plan B has an effectiveness of 0.60 + (0.40*0.75) = 0.9 or 90%

Vocaloid:

Select the correct answer. For what part of the population is Plan B NOT effective? A. 40% B. 30% C. 15% D. 10%